OpenStudy (anonymous):

Solve. 3√(m-1)-√(8m-1)=0

OpenStudy (asnaseer):

\[3\sqrt{m-1}-\sqrt{8m-1}=0\]\[3\sqrt{m-1}=\sqrt{8m-1}\]square both sides:\[9(m-1)=8m-1\]\[9m-9=8m-1\]\[9m-8m=9-1\]\[m=8\]